OpenStudy (anonymous):

Find the intercepts of the equation: 4x-6y=24

OpenStudy (anonymous):

re-arranging the equation we get: -6y=-4x+24 6y=4x-24 y=(4/6)x-(24/6) y=(2/3)-4 Slope of equation is (2/3) and intercept is 4 on negative y axis.

OpenStudy (anonymous):

y-int=-4 x-int=6

OpenStudy (anonymous):

My technique is setting the x value to 0 to solve y, and vice versa.

OpenStudy (anonymous):



OpenStudy (anonymous):